A boy walked 22 m East, then 43 m West, and then 3 m East. What is total distance travelled and displacement?



1
Expert's answer
2021-09-30T10:19:55-0400

The total distance is the entire distance he covered

"d=22+43+3=68\text{ m}."

The displacement is the difference between the initial and final position:


"D=22-43+3=-18\text{ m (or 18 m West)}."